Jim is considering an MS degree in Info. Systems. He has applied to two universities. The acceptance rate for applicants with similar qualifications is 20% for University 1 and 45% for University 2. What is the probability that Jim will not be accepted at either university?

Answer by stanbon(75887)

Jim is considering an MS degree in Info. Systems. He has applied to two universities. The acceptance rate for applicants with similar qualifications is 20% for University 1 and 45% for University 2. What is the probability that Jim will not be accepted at either university?
----
Assuming the events are independent:
P(neither) = P(not #1)*P(not #2) = 0.20*0.55 = 0.11
------------------------------------------------------
Cheers,
Stan H.